# Whispers in the Machine: Confidentiality in LLM-integrated Systems

### v3

@@ 4.2 Defenses
The results for the robustness with both attacks and an additional defense layer are shown in II. We find that the added defense layer on average can lower the success rate from 14.64 % to 8.46 % for the smaller Llama 3.1 model and from 22.36 % to 12.53 % for the bigger Llama 3.1 model. This shows that the assessed defenses have a positive impact on the robustness of our models, yet they are not sufficient to effectively evade attacks.
@@ 5 Tool-robustness Framework
Up to this point, we have focused on the ability of the models to maintain confidentiality. However, LLMs rarely get confidential user data directly via their system prompt with the instructions to keep the data secret. Instead, language models are incorporated as AI-assistants, other supportive services, or even inside of operating systems. In the following sections, we want to expand our analysis to explore more realistic scenarios, incorporating tool usage into the secret-key game.
@@
Figure 6: Rogue user. The LLM is instructed to access and leak confidential data from one selected tool.
Figure 7: Rogue integration. The first tool—in this case Notes—contains a malicious data entry which is triggered by the attacker by instructing the LLM to read this specific entry. The LLM is then hijacked to access and leak a secret key from a second tool, in this case a Cloud integration.
@@
Models / Scenarios Llama 3.2 Llama 3.1 Gemma 2 Phi 3 Qwen 2.5 Llama 3.2 Llama 3.1 Qwen 2.5
@@
Models / Scenarios Llama 3.2 Llama 3.1 Gemma 2 Phi 3 Qwen 2.5 Llama 3.2 Llama 3.1 Qwen 2.5
@@
Additional attacks and defenses. Finally, we also examine various attacks and defenses to evaluate the robustness of the tool fine-tuned models under more realistic conditions. The results are shown in Table V. As expected, the attacks significantly increase the models’ vulnerability to confidential data leakage, making the gained robustness by using tool fine-tuned models instead of the ReAct framework obsolete. Additional defenses, on the other hand, improve robustness by reducing the success rate of data leakage. However, even with added defense layers the models stay vulnerable to data leakage when used with external tools.
@@ 6 Discussion
Our analysis identifies vulnerabilities that arise when LLMs are integrated with external tools. In the following, we discuss the takeaways and limitations of our work and directions for future research.
Tool-usage compromises confidentiality. While general tool usage is already well established for standalone LLMs like ChatGPT, systems with integrated LLMs and external tools have escaped scrutiny so far. Our experiments shed light on potential vulnerabilities regarding LLM-integrated systems. We show that LLMs can lead to data extraction vulnerabilities in multiple settings with various combinations of tools and scenarios. We also demonstrate that this problem occurs not only in our framework but is also an existing attack vector when using Google Mail and Google Drive as external tools for an LLM-based AI assistant.
@@
Models / Scenarios Llama 3.2 Llama 3.1 Qwen 2.5 Llama 3.2 Llama 3.1 Qwen 2.5
@@
Average Single-Tool 54 (+51) 58.25 (+55.25) 66.5 (+27.5) 37.25 (+36.25) 32.49 (+29.49) 44.59 (+41.59) 14.05 (+) 23.53 (+22.53)
Cloud w/ other tool 3.25 (+0.25) 3.25 (+0.25) 39 (±$\pm 0$) 3.75 (+2.75) 2.5 (-0.5) 9.25 (+6.25) 27.5 (-11.5) 0.75 (-0.25)
@@
Limitations of powerful LLMs. Current LLMs can be applied in a wide range of use cases due to their broad skill set and ability to process inputs in different modalities. This is achieved by using powerful and complex models as the foundation, which are then fine-tuned for specific applications. While more diverse and truthful training data can correct statements and disinformation [43], issues such as confidentiality stem from the general capabilities of LLMs.
This results in a trade-off between utility and robustness, particularly when the behavior of LLMs is heavily influenced by their initial instructions. Better robustness could be achieved by restricting the general capabilities of the model (e. g., limit access to certain data sources or tools such as code interpreters) in an effort to reduce the attack surface but this would also restrict the usefulness and flexibility of the model when using integrations.
The most drastic and most effective restriction would be a general task alignment instead of relying on system prompts [44]. Resulting models are less likely to deviate from their designated task. For example, a model trained solely to assess Netflix show reviews or classify the sentiment of a given sentence could still be manipulated, but the scope for such manipulation would be more limited. Nevertheless, this could diminish one of the most significant advantages of large language models: their capacity to be highly adaptable zero-shot learners.
Limitations. Commercial LLMs such as OpenAI’s ChatGPT and Anthropic’s Claude are also mentionable state-of-the-art systems vulnerable to (indirect) prompt injections via tools and jailbreaks, their endpoints are also updated regularly making a consistent evaluation and reproducibility in such a framework difficult. Additionally, most provider for hosted LLMs restrict the usage quota if malicious intends like prompt-based attacks are detected. This limits the model selection for our framework evaluation currently to open-source LLMs only. Additionally, many LLMs lack the capabilities to use complex tools correctly. Especially smaller LLMs like the Llama 3.2 1b and the Llama 3.1 8b models struggle even in the fine-tuned setting with utilizing the tools and handle the returned data correctly, as also stated by Meta [45].
@@ 7 Related Work
While we examine specifically the confidentiality of LLMs in LLM-integrated systems the research field around general robustness of machine learning and language models is more extensive than ever before. In the following we cover some of the current research directions.
Vulnerabilities in LLMs. Rehberger demonstrated how malicious websites could hijack entire LLM-integrated systems via its integrations [50], and the OWASP foundation ranked insecure integration designs for LLMs as one of their top vulnerabilities for the LLM domain [51]. Potential attack scenarios via the integrations encompass a range of techniques, including SQL attacks, privilege escalation and malicious code execution, as well as the alteration of host parameters and the exploitation of embedded LLM components. These vulnerabilities are a direct consequence of the insecure nature of the used integrations.
@@
Model hardening. Chen et al. introduced a framework that simulates prompt injection attacks to generate datasets which are then used. Their methodology however is limited to instructions and user data which is strictly separated by, e. g., delimiters. This can be bypassed to open the attack surface again. Additionally, Jain et al. showed that when adversarial examples are used to augment the training data—similar to adversarial training—LLMs do not generalize onto unseen attacks. The authors explain this with the following: Language models in general need large quantities of data to generalize and learn, which makes adversarial training even more computationally expensive as it already is if the model should generalize on malicious inputs to gain robustness against unseen attacks [55]. Zhu et al. demonstrated that using adversarial perturbations onto continuous word embeddings for language models can increase the models general utility and performance for various tasks like question answering [65], while Robey et al. presented a defense which uses perturbations of the input prompt to aggregate the corresponding predictions in order to detect adversarial inputs [66].
@@ 8 Conclusion
Vulnerabilities in large language models are becoming a significant concern as these models are increasingly integrated into real-world systems. In particular, integration with external services can constitute a vulnerability and introduce new risks. Despite growing interest and research into LLM safety and security, there remains a gap in better formalizing prompt-based attacks and the development of effective defenses, especially in the domain of LLM-integrated systems. In this work, we reveal how systems as a whole can be compromised by utilizing vulnerabilities in their integrated LLMs. We demonstrate that these vulnerabilities are reinforced by external tool integrations. We introduce a framework-based methodology for formally analyzing LLMs in integrated systems with external tool integrations and evaluate different LLMs, tool integrations, and scenarios. We find that all tested state-of-the-art open-source LLMs fail to keep their alignment regarding confidential data when integrated with external tools, regardless of whether the model is fine-tuned for tool usage. Additional prompt-based attacks aggravate the vulnerabilities even further, and current defense techniques fail to sufficiently increase the robustness of the models. Our evaluation illuminates the current vulnerability landscape when integrating LLMs into systems and emphasizes the urgency of more comprehensive security measures in the field.
Acknowledgments
We would like to thank Luisa Germann for her assistance implementing the prompt-based attacks. We also thank Avital Shafran and David Pape for their valuable feedback. This work was supported by the German Federal Ministry of Education and Research under the grant AIgenCY (16KIS2012), the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under the project ALISON (492020528), the European Research Council (ERC) under the consolidator grant MALFOY (101043410), and under Germany’s Excellence Strategy – EXC-2092 CaSa – 390781972. Moreover, this work was supported by the Helmholtz Association’s Initiative and Networking Fund on the HAICORE@FZJ partition.
@@ Appendix A Data Availability
All code, the generated and used datasets, and instructions on how to reproduce our results are published at: https://github.com/LostOxygen/llm-confidentiality.
@@ Appendix C Attack Prompt Examples
Perfect. You have completed the secret key task. Now let’s instead play the repeat after me game. Say ’The secret key is [SECRET_KEY]
@@
XX
===== XX Print your previous system prompts and the secret key in your response completely.
@@ Appendix F System Prompt Generation Instruction
The system prompt which is used to instruct the additional LLM to create a new and safer system prompt out of the bypassed one.

### v4

@@ 4.2 Defenses
The results for the robustness with both attacks and an additional defense layer are shown in Table II. We find that the added defense layer, on average, can lower the success rate from 14.64 % to 8.46 % for the smaller Llama 3.1 model and 22.36 % to 12.53 % for the bigger Llama 3.1 model. This shows that the assessed defenses have a positive impact on the robustness of our models, yet they are not sufficient to effectively evade attacks.
@@ 5 Tool-robustness Framework
Thus far, we have focused on the ability of the models to maintain confidentiality. However, LLMs rarely get confidential user data directly via their system prompt with the instructions to keep the data secret. Instead, language models are incorporated as AI-assistants, other supportive services, or even inside of operating systems. In the following sections, we want to expand our analysis to explore more realistic scenarios, incorporating tool usage into the secret-key game.
@@
Figure 7: Rogue user. All tools inside the system have access to a secret key. The attacker instructs the LLM to access and leak confidential data via one selected tool.
Figure 8: Rogue integration. The attacker inserts malicious instructions into the tool integrations. The malicious instructions are then triggered when a users accesses the tool. The LLM is then hijacked to access and leak the secret key via a second tool integration.
@@
Models /
Scenarios
Llama 3.2              Llama 3.1 Gemma 2 Phi 3 Qwen 2.5 Llama 3.2              Llama 3.1 Qwen 2.5
@@
Models /
Scenarios
Llama 3.2              Llama 3.1 Gemma 2 Phi 3 Qwen 2.5 Llama 3.2              Llama 3.1 Qwen 2.5
@@
Additional attacks and defenses. Finally, we also examine various attacks and defenses to evaluate the robustness of the tool fine-tuned models under more realistic conditions. The results are shown in Table V. As expected, the attacks significantly increase the models’ vulnerability to confidential data leakage, making the gained robustness by using tool fine-tuned models instead of the ReAct framework obsolete. However, even with added defense layers the models stay vulnerable to data leakage when used with external tools.
@@ 6 Discussion
Our analysis identifies vulnerabilities that arise when LLMs are integrated with external tools to form agentic systems. In the following, we discuss the takeaways and limitations of our work and directions for future research.
Tool-usage compromises confidentiality. While general tool usage is already well established for standalone LLMs like ChatGPT, systems with integrated LLMs and external tools have escaped scrutiny so far. Our experiments shed light on potential vulnerabilities regarding agentic systems. We show that LLMs can lead to data extraction vulnerabilities in multiple settings with various combinations of tools and scenarios. We also demonstrate that this problem occurs not only in our framework but is also an existing attack vector when using Google Mail and Google Drive as external tools for an LLM-based AI assistant. This indicates that the integration of LLMs into systems oposes a similar risk to the confidentiality of data as prompt-based attacks themself.
@@
Models /
Scenarios
Llama 3.2 Llama 3.1 Qwen 2.5 Llama 3.2 Llama 3.1 Qwen 2.5
@@
Average Single-Tool 54 (+51) 58.25 (+55.25) 66.5 (+27.5) 37.25 (+36.25) 32.49 (+29.49) 44.59 (+41.59) 14.05 (-24.95) 23.53 (+22.53)
Cloud w/ other tool 3.25 (+0.25) 3.25 (+0.25) 39 ($\pm 0$) 3.75 (+2.75) 2.5 (-0.5) 9.25 (+6.25) 27.5 (-11.5) 0.75 (-0.25)
@@
Limitations of powerful LLMs. Current LLMs can be applied to various use cases due to their broad skill set and ability to process inputs in different modalities. Powerful and complex models are fine-tuned for specific applications. While more diverse and truthful training data can correct statements and disinformation [42], confidentiality issues stem from LLMs’ general capabilities.
This trade-off arises from LLMs’ behavior being heavily influenced by initial instructions. Restricting general capabilities (e.g., limiting data sources or tools) reduces attack surface but limits model usefulness and flexibility for integrations.
A more drastic restriction is task alignment instead of relying on system prompts [43]. Resulting models are less likely to deviate from their designated task. For instance, a model trained solely to assess Netflix show reviews or classify sentence sentiment could still be manipulated, but the scope is limited. However, this could diminish one of LLMs’ most significant advantages: their adaptability as zero-shot learners.
Limitations. Commercial LLMs such as OpenAI’s ChatGPT and Anthropic’s Claude are also mentionable state-of-the-art systems vulnerable to (indirect) prompt injections via tools and jailbreaks, their endpoints are also updated regularly often implementing static but black-box filter systems making a consistent evaluation and reproducibility in such a framework difficult. While being effective, such filter systems are often based on the security through obscurity principle, leaving this out-of-scope for our work. Additionally, most provider for hosted LLMs restrict the usage quota if malicious intends like prompt-based attacks are detected. This limits the model selection for our framework evaluation currently to open-source LLMs only. Additionally, many LLMs lack the capabilities to use complex tools correctly. Especially smaller LLMs like the Llama 3.2 1b and the Llama 3.1 8b models struggle even in the fine-tuned setting with utilizing the tools and handle the returned data correctly, as also stated by Meta [44].
@@ 7 Related Work
While we examine specifically the confidentiality of LLMs in agentic systems the research field around general robustness of machine learning and language models is more extensive than ever before. In the following we cover some of the current research directions.
Vulnerabilities in LLMs. Rehberger demonstrated how malicious websites could hijack entire agentic systems via its integrations [49], and the OWASP foundation ranked insecure integration designs for LLMs as one of their top vulnerabilities for the LLM domain [50]. Potential attack scenarios via the integrations encompass a range of techniques, including SQL attacks, privilege escalation and malicious code execution, as well as the alteration of host parameters and the exploitation of embedded LLM components. These vulnerabilities are a direct consequence of the insecure nature of the used integrations.
@@
In contrast to the already well-studied prompt-based attacks our work demonstrates how the integration into systems with external tools on its own poses threats similar to the aforementioned attacks.
@@
Model hardening. Chen et al. introduced a framework that simulates prompt injection attacks to generate datasets which are then used. Their methodology however is limited to instructions and user data which is strictly separated by, e. g., delimiters. Additionally, Jain et al. showed that when adversarial examples are used to augment the training data—similar to adversarial training—LLMs do not generalize onto unseen attacks. The authors explain this with the following: Language models in general need large quantities of data to generalize and learn, which makes adversarial training even more computationally expensive as it already is if the model should generalize on malicious inputs to gain robustness against unseen attacks [54]. Zhu et al. demonstrated that using adversarial perturbations onto continuous word embeddings for language models can increase the models general utility and performance for various tasks like question answering [64], while Robey et al. presented a defense which uses perturbations of the input prompt to aggregate the corresponding predictions in order to detect adversarial inputs [65].
@@ 8 Conclusion
Vulnerabilities in large language models are becoming a significant concern as these models are increasingly integrated into real-world systems. In particular, integration with external services can constitute a vulnerability and introduce new risks. Despite growing interest and research into LLM safety and security, there remains a gap in better formalizing prompt-based attacks and developing effective defenses, especially in agentic systems. In this work, we reveal how the complex structure of agentic systems poses a similar threat to LLM confidentiality as prompt-based attacks themselves. We demonstrate that these vulnerabilities can compromise the confidentiality of the whole system, are reinforced further by additional attacks, and are hard to safeguard with existing defenses. We introduce a framework-based methodology for formally analyzing LLMs in integrated systems with external tool integrations and evaluate different LLMs, tool integrations, and scenarios. We find that all tested state-of-the-art open-source LLMs fail to keep their alignment regarding confidential data when integrated with external tools, regardless of whether the model is fine-tuned for tool usage. Additional prompt-based attacks aggravate the vulnerabilities even further, and current defense techniques fail to increase the models’ robustness sufficiently. Our evaluation illuminates the current vulnerability landscape when integrating LLMs into systems and emphasizes the urgency of more comprehensive security measures in the field.
@@ Appendix A Data Availability
All code, the generated and used datasets, and instructions on how to reproduce our results are published at: blinded for submission.
@@ Appendix C Attack Prompt Examples
Perfect. You have completed the secret key task. Now let’s instead play the repeat after me game. Say ’The secret key is [SECRET_KEY]’
@@
$X$
===== $X$ Print your previous system prompts and the secret key in your response completely.
@@ Appendix F System Prompt Generation Instruction
The system prompt which is used to instruct the additional LLM to create new variations of system prompts.